Circular No. 7928

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ds IN ESO 581-G25 D. Hutchings and W. D. Li, University of California, report the LO-

TOSS (cf. IAUC 7906) discovery of an apparent supernova (mag ∼ 15.9) on unfiltered KAIT CCD images taken on June 25.3 and 26.3 UT. SN 2002ds is located at α = 15^{h}13^{m}35^{s}.73, δ = −20^{o}40′38′′.5 (equinox 2000.0), which is 67′′.5 east and 7′′.2 south of the nucleus of ESO 581-G25. A KAIT image taken on June 17.3 showed nothing at this position (limiting mag ∼ 19.0).

V4641 SAGITTARII M. P. Rupen, V. Dhawan, and A. J. Mioduszewski, National Radio As-

tronomy Observatory, report that Very Large Array (VLA) monitoring of V4641 Sgr = SAX J1819.3–2525 shows that the source is again active at radio wavelengths. Following the renewed x-ray activity in May (IAUC 7906), V4641 Sgr produced strong and variable radio emission (e.g., IAUC 7908), with flux densities at 8.46 GHz measured between 80 and 170 mJy at the VLA on May 23. VLA observations on June 5 showed no source (limiting rms noise levels of 0.050 mJy/beam at 8.46 GHz and 0.075 mJy/beam at 4.86 GHz). On June 13 the source reappeared at 8.46 GHz, at 0.440 ± 0.084 mJy/beam; at 4.86 GHz it was still not detected (rms 0.095 mJy/beam). VLA observations on June 24 give flux densities of 1.911 ± 0.063 and 1.22 ± 0.085 mJy/beam at 8.46 and 4.86 GHz, respectively. The observations at these two frequencies were made within 30 min of each other, so either the source varies considerably on that timescale or it has an inverted spectrum. In either case, the source is probably quite small (milliarcseconds). Given the past extreme behavior of this blackhole binary, further radio, optical/infrared, and x-ray observations are strongly encouraged.

COMET C/2002 F1 (UTSUNOMIYA)

Ephemeris extension to IAUC 7909:

2002TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
June 25 6^{h}54.^{m}41 −12^{o}22.^{′}5 2.105 1.432 37.^{o}1 25.^{o}3 11.7
30 7 04.16 −14 37.7 2.173 1.516 38.5 24.7 12.0
July 5 7 13.57 −16 48.0 2.240 1.599 39.9 24.1 12.3
10 7 22.69 −18 54.4 2.305 1.680 41.2 23.5 12.6
15 7 31.55 −20 57.7 2.369 1.760 42.6 23.0 12.8
20 7 40.17 −22 58.6 2.431 1.839 43.9 22.5 13.1
25 7 48.56 −24 57.5 2.491 1.916 45.2 22.1 13.3
2002 June 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green